General Meeting Summary 9/9/21
From Cuawiki
Jump to navigationJump to searchThe printable version is no longer supported and may have rendering errors. Please update your browser bookmarks and please use the default browser print function instead.
PARTICIPANTS: Aaron Arshak, Bogdan, Carlos, Hamlet, Paulo, Tanja, Peter, Emmanuel, Steven, Vardan, Vladimir, Julie, Marie, Rolf, Alex
HV DIVIDERS
- All dividers and PMTs have shipped from Ohio U. and have been received at JLab
- All tests have been completed from NPS collaboration
- Action item:
- define next steps, Fernando test etc.
LV CABLING STRATEGY
- Need to take a decision now to make progress
- Action item: schedule dedicated meeting with Fernando in the week of September 13th (before 11am ET) to decide on the path forward
- Participants: Bogdan, IJCLab-Orsay, Vladimir, ....
DSG UPDATES
- Graphical interface development EPICS temperature screens
- sensors are included at front and back
- Thermal analysis: 3x3 unit with with copper shell at constant temperature 10degC, added carbon fiber around individual blocks, applied heat of 0.5W from back
- temperature profile back to front looks very similar
- temperature profile lateral is different crystal to crystal - change between blocks: min~10, max~11.7degC
- Discussion why there is a change in temperature profile lateral (the outer crystals) from back to front
- There may be gaps between crystals in the model - this would effectively make the inner crystal thermally isolated -- need to check
- Note: important for the experiment is the temperature variation crystal to crystal - don't want different light yield crystal to crystal
- light yield change 2%/degC
- ideally want sub-degree variation
- Database being created of all NPS plots, e.g., CAEN model tests, HV supply tests, etc.
- Actions:
- Check on air gaps in the model
- Thermal analysi of the full (36x31) NPS assembly
PRESHAPING ACTIVITY
- 2-3 people from DSG will participate
- currently have available 4 jigs - will produce 6 more
- ODU tech will help too
- provided 6 jigs and 500 foils
- start with 30-40 foils then quality check
- Action item:
- work our arrangement for pre-shaping activity space at JLab
- Produce additional 6 jigs
- Check in on ODU progress
MECHANICAL
- Work ongoing, e.g., detector lift fixture, cables, etc.
CRYSTALS
- Have been receiving ~40 crystals/month from CRYTUR
- all good quality
- most of them now with the "new" raw material (available from vendor instead of using the raw material from Giessen U.)
NEXT MEETING: THURSDAY 30 SEPTEMBER AT 9AM ET